# Magnitude homology of real hyperplane arrangements

### 04 Apr 2026 v1

@@ 2.6. Varchenko matrix
Varchenko introduced a bilinear form and computed the determinant [31]. Generalization to oriented matroids can be found in [17]. See [8] and Sections 8.4-8.5 of [2] for details.
Here we only consider the equal weight case. For a real hyperplane arrangement $\mathcal{A}$, we define the $q$-Varchenko matrix
@@
In other words, $\mathcal{V}_{q}(\mathcal{A})$ is the Zeta matrix of the tope graph $\mathcal{T}(\mathcal{A})$ (see Section 2.2). It can also be defined for oriented matroids.
Theorem 2.33 ([31], see also Theorem 8.23 of [2]).
@@
$\det\mathcal{V}_{q}(\mathcal{A})=\prod_{X\in L(\mathcal{A})\setminus\{\hat{0}\}}(1-q^{2|\mathcal{A}_{X}|})^{c^{X}\beta_{X}}$
@@ 2.7. Tope poset and generating function
Fix a base chamber $B\in\mathrm{Ch}(\mathcal{A})$ and define a partial order on $\mathrm{Ch}(\mathcal{A})$ by $C\preceq C^{\prime}$ if $S(B,C)\subseteq S(B,C^{\prime})$. This poset is called the tope poset of $\mathcal{A}$ based at $B$, denoted by $\mathcal{T}(\mathcal{A},B)$. The poset $\mathcal{T}(\mathcal{A},B)$ has bottom element $B$ and top element B-B, graded by rank function $\rho_{B}(C)=d(B,C)$. In terms of notations in Section 2.1, $\mathcal{T}(\mathcal{A},B)=[B,-B]_{\mathcal{T}(\mathcal{A})}$.
Proposition 2.34 (Theorem 2.2 of [10], see also Theorem 4.4.2 of [4]).
An open interval $(C_{1},C_{2})$ inside $\mathcal{T}(\mathcal{A},B)$ is homotopy equivalent to a sphere $S^{\mathrm{rank}\mathcal{A}-\dim F-2}$, if the closed interval $C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$; contractible otherwise.
Corollary 2.35 (Theorem 1.11 of [11], see also Corollary 4.4.3 of [4]).
The Möbius function on $\mathcal{T}(\mathcal{A},B)$ is
$\mu(C_{1},C_{2})=\begin{cases}(-1)^{\mathrm{rank}\mathcal{A}-\dim F},&\text{ if }[C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F},\exists F\in\mathcal{F}(\mathcal{A});\\ 0,&\text{ otherwise}.\end{cases}$
Proposition 2.36 (Theorem 3.4 of [3]).
@@
Let $D_{\mathcal{A},B}(q)=\sum_{C\in\mathrm{Ch(\mathcal{A})}}q^{d(B,C)}$ be the rank generation function of $\mathcal{T}(\mathcal{A},B)$. Note that it is the sum of entries of the $B$-row/column of the $q$-Varchenko matrix $\mathcal{V}_{q}(\mathcal{A})$.
@@
If $\mathcal{A}$ is a Coxeter arrangement with exponents $e_{1},\ldots,e_{\ell}$, then
@@
for every $B\in\mathrm{Ch}(\mathcal{A})$, where $n]_{q}=1+q+q^{2}+\cdots+q^{n-1}$ is the $q$-analog.
The generating function decomposition formula is generalized to supersolvable arrangements [3], inductively factored arrangements [18]. See also [1, 26, 12].
3. Magnitude of tope graphs
For a real hyperplane arrangement $\mathcal{A}$, we define the magnitude $\mathrm{Mag}(\mathcal{A})=\mathrm{Mag}(\mathcal{A},q)$ of $\mathcal{A}$ as the magnitude of $\mathcal{T}(\mathcal{A})$.
@@ 3.1. Basic properties
We first prove a reciprocity result.
@@
Let $\mathcal{A}$ be a central hyperplane arrangement with $N=\#\mathcal{A}$ hyperplanes. Then
@@
Recall that there is an involution $\tau:\mathrm{Ch}(\mathcal{A})\to\mathrm{Ch}(\mathcal{A})$ sending $D$ to D-D. For any two chambers $C,D\in\mathrm{Ch}(\mathcal{A})$, we have d(C,D)+d(C,D)=Nd(C,D)+d(C,-D)=N. Then the $q$-Varchenko matrix $\mathcal{V}_{q}=\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]$ satisfies
@@
where $P_{\tau}$ is the permutation matrix induced by $\tau$. By definition of magnitude,
@@
Since tope graphs are partial cubes, the magnitude $\mathrm{Mag}(\mathcal{A})$ has the one-point property by [22].
@@
Let $\mathcal{A}$ be a central hyperplane arrangement, then $\mathrm{Mag}(\mathcal{A},1)=1$.
Next we prove some main structural results of $\mathrm{Mag}(\mathcal{A})$.
@@
For central arrangement $\mathcal{A}$, let $\mathrm{Mag}(\mathcal{A},q)=P(q)/Q(q)$ be the reduced form, with the leading coefficient of $Q(q)$ positive. Then
@@
    Roots of $Q$ are all roots of unity, excluding $1$.
@@
Let $P(q)=a_{m}q^{m}+\cdots+a_{0}$ and $Q(q)=b_{n}q^{n}+\cdots+b_{0}$ where $a_{m}\neq 0$ and $b_{n}\neq 0$. Then $m=\deg P$ and $n=\deg Q$. Since $\mathrm{Mag}(\mathcal{A},0)=a_{0}/b_{0}=\#\mathrm{Ch}(\mathcal{A})$ (Proposition 2.7), we have $a_{0}\neq 0$ and $b_{0}\neq 0$. By the reciprocity formula (Proposition 3.1),
@@
Multiplying both sides by $q^{m-n}$, we have
@@
Comparing degrees, we conclude N+mn=0N+m-n=0 and (1) is proved.
Let us write $P^{*}(q)=q^{m}P(q^{-1})=a_{m}+\cdots+a_{0}q^{m}$ and $Q^{*}(q)=q^{n}Q(q^{-1})=b_{n}+\cdots+b_{0}q^{n}$ as the reverse polynomials. We have proved
@@
Note that the fraction $P^{*}(q)/Q^{*}(q)$ is reduced since if $P^{*}$ and $Q^{*}$ share a common root $\alpha\neq 0$, $P$ and $Q$ would share a common root $1/\alpha$, contradicting to the assumption $P/Q$ is reduced. Now we must have $P^{*}(q)=rP(q)$ and $Q^{*}(q)=rQ(q)$ for some constant $r\neq 0$. Substituting $q=1$ into the equations $Q^{*}(q)=q^{n}Q(1/q)$ and $Q^{*}(q)=rQ(q)$, we obtain
@@
Since $\mathrm{Mag}(\mathcal{A},1)=P(1)/Q(1)=1$ (Proposition 3.2) and $P/Q$ is reduced, $Q(1)\neq 0$, we conclude $r=1$. Then $P^{*}=P$ and $Q^{*}=Q$, so (2) is proved.
Since $\mathcal{V}_{q}^{-1}=\frac{1}{\det\mathcal{V}_{q}}\mathrm{adj~}\mathcal{V}_{q}$, roots of $Q(q)$ are all roots of $\det\mathcal{V}_{q}$, which are roots of unity by Theorem 2.33. We have already seen $Q(1)\neq 0$. So (3) is proved. ∎
If an arrangement $\mathcal{A}$ decomposes as a direct sum $\mathcal{A}_{1}\oplus\mathcal{A}_{2}$, then the tope graph decomposes as cartesian product $\mathcal{T}(\mathcal{A})=\mathcal{T}(\mathcal{A}_{1})\square\mathcal{T}(\mathcal{A}_{2})$. So by Proposition 2.5, we have a decomposition of magnitude.
@@
Coxeter arrangements have the natural action of the Coxeter groups on the chambers, which makes the magnitude easy to compute.
@@
For Coxeter arrangement $\mathcal{A}_{W}$ with exponents $\{e_{1},\ldots,e_{\ell}\}$, we have
@@
The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is the $1$-skeleton of the $W$-permutohedron, or equivalently the Cayley graph of $W$ with respect to the standard generating set (Example 2.29), hence is vertex-transitive. The result follows from Proposition 2.3 and Proposition 2.37. ∎
3.2. Face decomposition conjecture
For the purpose of reciprocity, let us define the following.
@@ 3.2. Face decomposition conjecture
$\mathrm{Mag}^{\circ}(\mathcal{A},q):=(-1)^{\mathrm{rank}\mathcal{A}}\mathrm{Mag}(\mathcal{A},q^{-1})=(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}\mathrm{Mag}(\mathcal{A},q),$
where the latter equality is Proposition 3.1.
Let us propose a main conjecture of the magnitude of arrangements.
Conjecture 3.7.
We have a face decomposition of the magnitude.
$\mathrm{Mag}(\mathcal{A})$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\mathrm{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})$
$=\sum_{X\in L(\mathcal{A})}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})=\sum_{X\in L(\mathcal{A})}(-1)^{\mathrm{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X}),$
where $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$. Equivalently,
$\mathrm{Mag}(\mathcal{A})$ $=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\mathrm{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}$
$=\frac{\sum_{X\in L(\mathcal{A})\setminus\{\cap\mathcal{A}\}}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{X\in L(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\mathrm{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X})}{1-(-1)^{\mathrm{rank}\mathcal{A}}q^{\#\mathcal{A}}}.$
@@
This formula is the discrete $q$-analogue of the decomposition of the Euler characteristic of a zonotope into its open faces. Conjecture 3.7 is verified by computer for all examples listed in Section 3.3 and the appendix, and by easy computations for Boolean arrangements and all rank $2$ arrangements.
@@
Conjecture 3.7 implies that $\mathrm{Mag}(\mathcal{A})$ is determined by $L(\mathcal{A})$.

### 01 May 2026 v3

@@ 2.6. Varchenko matrix
Varchenko introduced a bilinear form associated with an arrangement and computed its determinant [33]. Generalizations to oriented matroids can be found in [18]. See [8] and [2, Sections 8.4–8.5] for details.
Here we only consider the equal-weight case. For a real hyperplane arrangement $\mathcal{A}$, we define the $q$-Varchenko matrix by
@@
In other words, $\mathcal{V}_{q}(\mathcal{A})$ is the Zeta matrix of the tope graph $\mathcal{T}(\mathcal{A})$ (see Section 2.2). This matrix can also be defined for oriented matroids.
Theorem 2.34 ([33], see also [2, Theorem 8.23]).
@@
$\det\mathcal{V}_{q}(\mathcal{A})=\prod_{X\in L(\mathcal{A})\setminus\{\hat{0}\}}(1-q^{2|\mathcal{A}_{X}|})^{c^{X}\beta_{X}}.$
@@ 2.7. Tope poset and generating function
Fix a chamber $B\in\mathrm{Ch}(\mathcal{A})$ and define a partial order on $\mathrm{Ch}(\mathcal{A})$ by $C\preceq C^{\prime}$ if $S(B,C)\subseteq S(B,C^{\prime})$. This poset is called the tope poset of $\mathcal{A}$ based at $B$, denoted by $\mathcal{T}(\mathcal{A},B)$. The poset $\mathcal{T}(\mathcal{A},B)$ has bottom element $B$ and top element B-B, graded by the rank function $\rho_{B}(C)=d(B,C)$. In the notation of Section 2.1, $\mathcal{T}(\mathcal{A},B)=[B,-B]_{\mathcal{T}(\mathcal{A})}$.
Proposition 2.35 ([9, Theorem 2.2], see also [4, Theorem 4.4.2]).
An open interval $(C_{1},C_{2})$ in $\mathcal{T}(\mathcal{A},B)$ is homotopy equivalent to a sphere $S^{\operatorname{rank}\mathcal{A}-\dim F-2}$ if $C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F}$ for some $F\in\mathcal{F}(\mathcal{A})$, and is contractible otherwise.
Corollary 2.36 ([10, Theorem 1.11], see also [4, Corollary 4.4.3]).
The Möbius function on $\mathcal{T}(\mathcal{A},B)$ is given by
$\mu(C_{1},C_{2})=\begin{cases}(-1)^{\operatorname{rank}\mathcal{A}-\dim F},&\text{ if }[C_{1},C_{2}]=\mathrm{Ch}(\mathcal{A})_{F}\text{ for some }F\in\mathcal{F}(\mathcal{A});\\ 0,&\text{ otherwise}.\end{cases}$
Proposition 2.37 ([3, Theorem 3.4]).
@@
Let $D_{\mathcal{A},B}(q)=\sum_{C\in\mathrm{Ch}(\mathcal{A})}q^{d(B,C)}$ be the rank-generating function of $\mathcal{T}(\mathcal{A},B)$. Note that it is the sum of the entries in the $B$-row/column of the $q$-Varchenko matrix $\mathcal{V}_{q}(\mathcal{A})$.
@@
If $\mathcal{A}$ is a Coxeter arrangement with exponents $e_{1},\ldots,e_{\ell}$, then we have
@@
for every $B\in\mathrm{Ch}(\mathcal{A})$, where $n]_{q}=1+q+q^{2}+\cdots+q^{n-1}$ is the $q$-number.
The generating-function decomposition formula has been generalized to supersolvable arrangements [3] and inductively factored arrangements [19]. See also [1, 27, 12].
3. Magnitude of arrangements
For a real central hyperplane arrangement $\mathcal{A}$, we define the magnitude $\mathrm{Mag}(\mathcal{A})=\mathrm{Mag}(\mathcal{A},q)$ of $\mathcal{A}$ as the magnitude of its tope graph $\mathcal{T}(\mathcal{A})$.
@@ 3.1. Basic properties
We first prove the symmetry of the magnitude with respect to $q\leftrightarrow q^{-1}$.
@@
Let $\mathcal{A}$ be an arrangement with $N=\#\mathcal{A}$ hyperplanes. Then
@@
Recall that there is an involution $\tau:\mathrm{Ch}(\mathcal{A})\to\mathrm{Ch}(\mathcal{A})$ sending $D$ to D-D. For any two chambers $C,D\in\mathrm{Ch}(\mathcal{A})$, we have d(C,D)+d(C,D)=Nd(C,D)+d(C,-D)=N. Thus the $q$-Varchenko matrix $\mathcal{V}_{q}=\mathcal{V}_{q}(\mathcal{A})=[q^{d(C,D)}]$ satisfies
@@
where $P_{\tau}$ is the permutation matrix induced by $\tau$. By the definition of magnitude, we have
@@
Since tope graphs are partial cubes, the magnitude $\mathrm{Mag}(\mathcal{A})$ satisfies the one-point property established in [22].
@@
If $\mathcal{A}$ is an arrangement, then $\mathrm{Mag}(\mathcal{A},1)=1$.
Next we prove some structural results for $\mathrm{Mag}(\mathcal{A})$.
@@
For an arrangement $\mathcal{A}$, let $\mathrm{Mag}(\mathcal{A},q)=P(q)/Q(q)$ be the reduced expression of the magnitude, with the leading coefficient of $Q(q)$ positive. Then
@@
    All roots of $Q$ are roots of unity other than $1$.
@@
Let $P(q)=a_{m}q^{m}+\cdots+a_{0}$ and $Q(q)=b_{n}q^{n}+\cdots+b_{0}$ where $a_{m}\neq 0$ and $b_{n}\neq 0$. Then $m=\deg P$ and $n=\deg Q$. Since $\mathrm{Mag}(\mathcal{A},0)=a_{0}/b_{0}=\#\mathrm{Ch}(\mathcal{A})$ (Proposition 2.7), we have $a_{0}\neq 0$ and $b_{0}\neq 0$. By the symmetry formula (Proposition 3.1), we have
@@
Multiplying both sides by $q^{m-n}$, we obtain
@@
Comparing degrees, we obtain N+mn=0N+m-n=0, and (1) is proved.
Write $P^{*}(q)=q^{m}P(q^{-1})=a_{m}+\cdots+a_{0}q^{m}$ and $Q^{*}(q)=q^{n}Q(q^{-1})=b_{n}+\cdots+b_{0}q^{n}$ for the reverse polynomials. We have proved
@@
Note that the fraction $P^{*}(q)/Q^{*}(q)$ is reduced, since if $P^{*}$ and $Q^{*}$ share a common root $\alpha\neq 0$, then $P$ and $Q$ would share the common root $1/\alpha$, contradicting the assumption that $P/Q$ is reduced. Hence we must have $P^{*}(q)=rP(q)$ and $Q^{*}(q)=rQ(q)$ for some constant $r\neq 0$. Substituting $q=1$ into the equations $Q^{*}(q)=q^{n}Q(1/q)$ and $Q^{*}(q)=rQ(q)$, we obtain
@@
Since $\mathrm{Mag}(\mathcal{A},1)=P(1)/Q(1)=1$ (Proposition 3.2) and $P/Q$ is reduced, we have $Q(1)\neq 0$, so $r=1$. Then $P^{*}=P$ and $Q^{*}=Q$, and (2) is proved.
Since $\mathcal{V}_{q}^{-1}=\frac{1}{\det\mathcal{V}_{q}}\mathrm{adj~}\mathcal{V}_{q}$, the roots of $Q(q)$ are roots of $\det\mathcal{V}_{q}$, which are roots of unity by Theorem 2.34. We have already seen that $Q(1)\neq 0$. Thus (3) is proved. ∎
If an arrangement $\mathcal{A}$ decomposes as a direct sum $\mathcal{A}_{1}\oplus\mathcal{A}_{2}$, then the tope graph decomposes as the Cartesian product $\mathcal{T}(\mathcal{A})=\mathcal{T}(\mathcal{A}_{1})\square\mathcal{T}(\mathcal{A}_{2})$. Therefore, by Proposition 2.5, we have a decomposition of magnitude.
@@
Coxeter arrangements carry a natural action of the corresponding Coxeter groups on the chambers, which makes their magnitudes easy to compute.
@@
For a Coxeter arrangement $\mathcal{A}_{W}$ with exponents $\{e_{1},\ldots,e_{\ell}\}$, we have
@@
The tope graph $\mathcal{T}(\mathcal{A}_{W})$ is the $1$-skeleton of the $W$-permutahedron, or equivalently the Cayley graph of $W$ with respect to the standard generating set (Example 2.32), and hence is vertex-transitive. The result follows from Proposition 2.3 and Proposition 2.38. ∎
3.2. Face decomposition formula
For reciprocity purposes, let us define the following.
@@ 3.2. Face decomposition formula
$\mathrm{Mag}^{\circ}(\mathcal{A},q):=(-1)^{\operatorname{rank}\mathcal{A}}\mathrm{Mag}(\mathcal{A},q^{-1})=(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}\mathrm{Mag}(\mathcal{A},q),$
where the latter equality follows from Proposition 3.1.
We prove the following “face decomposition formula” for the magnitude of an arrangement.
Theorem 3.7.
For any arrangement $\mathcal{A}$, we have
$\mathrm{Mag}(\mathcal{A})=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F}).$
@@
Theorem 3.7 can be reformulated as
$\mathrm{Mag}(\mathcal{A})=\sum_{X\in L(\mathcal{A})}c^{X}\mathrm{Mag}^{\circ}(\mathcal{A}_{X})=\sum_{X\in L(\mathcal{A})}(-1)^{\operatorname{codim}X}c^{X}q^{\#\mathcal{A}_{X}}\mathrm{Mag}(\mathcal{A}_{X}),$
where $c^{X}=\#\mathrm{Ch}(\mathcal{A}^{X})$. It can also be written as
$\mathrm{Mag}(\mathcal{A})=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}\mathrm{Mag}^{\circ}(\mathcal{A}_{F})}{1-(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}}=\frac{\sum_{F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F})}{1-(-1)^{\operatorname{rank}\mathcal{A}}q^{\#\mathcal{A}}}.$
This can be regarded as a recursive formula for computing $\mathrm{Mag}(\mathcal{A})$.
Remark 3.9.
Taking the limit $q\to 1$ gives $\mathrm{Mag}(\mathcal{A},1)=1$ and $\mathrm{Mag}^{\circ}(\mathcal{A},1)=(-1)^{\operatorname{rank}\mathcal{A}}$ by Proposition 3.2. Therefore, Theorem 3.7 can be regarded as a $q$-analogue of the decomposition of the Euler characteristic of the zonotope into its open faces.
@@
The magnitude $\mathrm{Mag}(\mathcal{A})$ of an arrangement $\mathcal{A}$ is determined by $L(\mathcal{A})$.
Lemma 3.11.
For any arrangement $\mathcal{A}$ and $D\in\mathrm{Ch}(\mathcal{A})$, we have
$\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}+d(D,FD)}=1.$
Proof.
If $\operatorname{rank}\mathcal{A}=0$, then $\mathcal{A}$ is the empty arrangement and there is nothing to prove, so let us assume $r:=\operatorname{rank}\mathcal{A}\geq 1$. Passing to the essentialization does not change the chamber set, the tope graph, or the quantities $\mathcal{A}_{F}$ and $\operatorname{codim}F=\operatorname{rank}(\mathcal{A}_{F})$, so we may work in an essential arrangement of rank $r$.
Choose defining forms $\alpha_{H}$ so that $D_{H}=+$ for every $H\in\mathcal{A}$. For a face $F$, define $M_{D}(F)\subset\mathcal{A}$ by
$M_{D}(F):=\{H\in\mathcal{A}:F_{H}\neq+\}.$
Since $F_{H}=0$ exactly when $H\in\mathcal{A}_{F}$, while $d(D,FD)$ counts the hyperplanes on which $F_{H}=$-, we have $M_{D}(F)=\#\mathcal{A}_{F}+d(D,FD)$. Therefore it is enough to prove
$S_{D}(q):=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#M_{D}(F)}=1.$
Expand each power as
$q^{\#M_{D}(F)}=(1+(q-1))^{\#M_{D}(F)}=\sum_{T\subseteq M_{D}(F)}(q-1)^{\#T}.$
After exchanging the order of summation we obtain
$S_{D}(q)=\sum_{T\subseteq\mathcal{A}}(q-1)^{\#T}E_{T},\qquad E_{T}:=\sum_{F:\,T\subseteq M_{D}(F)}(-1)^{\operatorname{codim}F}.$
Hence the desired identity will follow once we show $E_{T}=\mathbf{1}_{T=\varnothing}$, i.e. $E_{T}=0$ if $T\neq\varnothing$ and $E_{\varnothing}=1$. We interpret $E_{T}$ as a reduced Euler characteristic. Let $S^{r-1}$ be the unit sphere in the ambient space $V=\mathbb{R}^{r}$. For each hyperplane $H\in\mathcal{A}$, let
$H_{D}^{-}:=\{x\in S^{r-1}:\alpha_{H}(x)\leq 0\}$
be the closed hemisphere opposite to $D$, and for $T\subseteq\mathcal{A}$ define
$\Sigma_{T}:=\bigcap_{H\in T}H_{D}^{-}\subseteq S^{r-1},\qquad\Sigma_{\varnothing}:=S^{r-1}.$
The arrangement induces a regular CW decomposition of $S^{r-1}$ where the open cells are $F\cap S^{r-1}$ for non-minimal faces $F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}$. An open cell $F\cap S^{r-1}$ lies in $\Sigma_{T}$ if and only if $T\subseteq M_{D}(F)$. Therefore
$\chi(\Sigma_{T})=\sum_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}\\ T\subseteq M_{D}(F)\end{subarray}}(-1)^{\dim F-1}.$
Since $\operatorname{codim}F=r-\dim F$ and the minimal face $\cap\mathcal{A}$ contributes $(-1)^{r}$ to $E_{T}$, we obtain
$E_{T}$ $=(-1)^{r}+\sum_{\begin{subarray}{c}F\in\mathcal{F}(\mathcal{A})\setminus\{\cap\mathcal{A}\}\\ T\subseteq M_{D}(F)\end{subarray}}(-1)^{r-\dim F}=(-1)^{r-1}\left(\chi(\Sigma_{T})-1\right).$
If $T=\varnothing$, then $\Sigma_{\varnothing}=S^{r-1}$, so $E_{\varnothing}=(-1)^{r-1}(\chi(S^{r-1})-1)=1$. If $T\neq\varnothing$, then $\Sigma_{T}$ is the intersection of a nonempty collection of closed hemispheres, so $\Sigma_{T}$ is homeomorphic to $D^{r-1}$. Therefore $E_{T}=(-1)^{r-1}(\chi(D^{r-1})-1)=0$. This shows $E_{T}=\mathbf{1}_{T=\varnothing}$. ∎
Proof of Theorem 3.7.
Fix a face $F\in\mathcal{F}(\mathcal{A})$. Let $w_{F}\colon\mathrm{Ch}(\mathcal{A})_{F}\to\mathbb{Q}(q)$ be the weight vector of $\mathcal{T}(\mathcal{A})_{F}\cong\mathcal{T}(\mathcal{A}_{F})$, so that
$\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}q^{d(E,C)}w_{F}(C)=1$
for every $E\in\mathrm{Ch}(\mathcal{A})_{F}$. Extend $w_{F}$ by zero outside $\mathrm{Ch}(\mathcal{A})_{F}$. Now define a function $\widetilde{w}_{\mathcal{A}}\colon\mathrm{Ch}(\mathcal{A})\to\mathbb{Q}(q)$ by
$\widetilde{w}_{\mathcal{A}}(C):=\sum_{F\leq C}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}w_{F}(C).$
Taking the sum over all chambers gives
$\sum_{C\in\mathrm{Ch}(\mathcal{A})}\widetilde{w}_{\mathcal{A}}(C)$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}w_{F}(C)$
$=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\mathrm{Mag}(\mathcal{A}_{F},q)$
$=\sum_{F\in\mathcal{F}(\mathcal{A})}\mathrm{Mag}^{\circ}(\mathcal{A}_{F},q).$
Therefore Theorem 3.7 will follow once we prove that $\widetilde{w}_{\mathcal{A}}$ is the actual weight vector of $\mathcal{A}$. Fix a chamber $D\in\mathrm{Ch}(\mathcal{A})$. We compute the $D$-th coordinate of $\mathcal{V}_{q}(\mathcal{A})\widetilde{w}_{\mathcal{A}}$:
$\sum_{C\in\mathrm{Ch}(\mathcal{A})}q^{d(D,C)}\widetilde{w}_{\mathcal{A}}(C)$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}}\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}q^{d(D,C)}w_{F}(C).$
Since $\mathcal{T}(\mathcal{A})_{F}$ is gated with gate projection $D\mapsto FD$ (Lemma 2.29), we have
$d(D,C)=d(D,FD)+d(FD,C)$
for every $C\in\mathrm{Ch}(\mathcal{A})_{F}$. Using this identity and the defining equation for the local weight vector $w_{F}$, we obtain
$\sum_{C\in\mathrm{Ch}(\mathcal{A})}q^{d(D,C)}\widetilde{w}_{\mathcal{A}}(C)$ $=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}+d(D,FD)}\sum_{C\in\mathrm{Ch}(\mathcal{A})_{F}}q^{d(FD,C)}w_{F}(C)$
$=\sum_{F\in\mathcal{F}(\mathcal{A})}(-1)^{\operatorname{codim}F}q^{\#\mathcal{A}_{F}+d(D,FD)}.$
By Lemma 3.11, this equals $1$ for every $D\in\mathrm{Ch}(\mathcal{A})$. ∎